# Predictability Enables Parallelization of Nonlinear State Space Models

### v1

@@ 6 Conclusion
The DeepPCR algorithm was introduced by Danieli et al. (2023), who investigated its convergence rates empirically but not theoretically. Around the same time, Lim et al. (2024) independently proposed an essentially identical method under the name DEER, proving local quadratic convergence for Gauss–Newton under standard assumptions but leaving the question of global convergence unresolved. Gonzalez et al. (2024) proved the global convergence of DEER and other variants, though only with worst-case bounds of $T$ optimization steps. None of these prior works addressed the relationship between system dynamics and conditioning, or established global linear convergence rates.
@@
Global convergence rates for Gauss-Newton are rare, despite the breadth of optimization literature (Nocedal and Wright, 2006; Boyd and Vandenberghe, 2004; Nesterov, 2018). Theorem 4 establishes global convergence with linear rate for Gauss-Newton by leveraging our specific problem structure.
@@
Fifty years ago, Hyafil and Kung (1975) and Kung (1976) showed that linear recursions enjoy speedups from parallel processors while nonlinear recursions of rational functions with degree larger than one cannot. These prescient works set the stage for our more general findings, which explicitly link the dynamical properties of the recursion to its parallelizability. Parallel-in-time methods for continuous systems also have a long history (Gander, 2015; Ong and Schroder, 2020), with Chartier and Philippe (1993) showing that dissipative systems can be parallelized using multiple shooting. Furthermore, Danieli and MacLachlan (2021) and De Sterck et al. (2025) identify the CFL number as an important system quantity for determining the usefulness of multigrid systems; drawing a deeper connection between this line of work and our paper is an interesting direction for future research.
@@
More recently, several works have parallelized diffusion models via fixed-point iteration (Danieli et al., 2023; Shih et al., 2023; Tang et al., 2024; Selvam et al., 2024), again with $T$-step worst-case guarantees. Anari et al. (2024) proved $\log(T)$ rates for a particular dynamical system (Langevin dynamics) and a particular fixed-point iteration (Picard iteration). Crucially, prior work has not focused on the merit function, which we can define for any discrete-time dynamical system and optimizer.
@@
To our knowledge, no prior work connects the LLE of a dynamical system to the conditioning of the corresponding optimization landscape, as established in Theorem˜2. In particular, we showed that systems with high unpredictability yield poorly conditioned (i.e., flat) merit functions, linking dynamical instability to optimization difficulty in a geometrically appealing way.
@@
The centrality of parallel sequence modeling architectures like transformers (Vaswani et al., 2017), deep SSMs (Gu et al., 2021; Smith et al., 2023; Gu and Dao, 2023), and linear RNNs (Yang et al., 2024) in modern machine learning underscores the need for our theoretical work. Merrill et al. (2024) explored the question of parallelizability through the lens of circuit complexity, analyzing when dynamical-system-based models can solve structured tasks in constant depth. Their focus complements ours, and suggests an opportunity for synthesis in future work.
@@
Nonlinear state space models are ubiquitous across science, engineering, and machine learning. Our results offer two key contributions.
@@
First, they provide a principled way to determine, a priori, whether optimization-based parallelization of a given model is practical. In many physical, robotic, and control systems, particularly those that are strongly dissipative, this insight enables orders-of-magnitude speed-ups on GPU hardware (Kolter and Manek, 2019; Beik-Mohammadi et al., 2024; Jaffe et al., 2024; Fan et al., 2022; Sindhwani et al., 2018; Sun et al., 2021; Tsukamoto et al., 2021; Revay et al., 2023).
@@
In concurrent work, Zoltowski et al. (2025) developed and leveraged quasi-Newton methods to parallelize Markov Chain Monte Carlo over the sequence length, attaining order of magnitude speed-ups in wallclock time. These speed-ups were made possible by the fast convergence of the quasi-Newton methods in the settings considered by Zoltowski et al. (2025). Suggestively, there is an abundance of research studying the contractivity of MCMC in different settings (Bou-Rabee et al., 2020; Mangoubi and Smith, 2021). In fact, the empirical results in this paper showing that Langevin dynamics can have negative LLE (cf. Figure˜3) are suggestive that the Metropolis-adjusted Langevin algorithm (MALA), a workhorse of MCMC, may also be predictable in settings of interest. A precise characterization of what makes an MCMC algorithm and target distribution contractive (i.e. predictable, in the language of our paper) would provide useful guidance for when one should aim to parallelize MCMC over the sequence length. Thus, providing precise theoretical justification for parallelizing MCMC over the sequence length is an exciting avenue for future work.
@@
Second, our results have direct implications for system design. When constructing nonlinear dynamical systems in machine learning—such as novel recurrent neural networks—parallelization benefits are maximized when the system is deliberately made predictable. Given the large body of work on training stable RNNs (Miller and Hardt, 2019; Erichson et al., 2020; Kozachkov et al., 2022; Goel et al., 2022; Krotov, 2023; Engelken, 2023; Revay et al., 2023; Orvieto et al., 2023; Jaffe et al., 2024; Farsang et al., 2025), many effective techniques already exist for enforcing stability or predictability during training. A common approach is to parameterize the model’s weights so that all intermediate models encountered during training are predictable by construction, thereby guaranteeing that the final trained model is predictable as well. We provide a simple parameterization of a contractive SSM in Appendix˜I.
@@
Notably, the concurrent work of Farsang et al. (2025) develops a nonlinear SSM and trains it in parallel using DEER. Farsang et al. (2025) explicitly parameterizes their LrcSSM to be contractive (see their Appendix A.1). While they do not identify it as such, this contractivity is precisely what enables DEER to converge quickly throughout training. Ensuring a negative largest Lyapunov exponent through parameterization guarantees parallelizability for the entire training process, enabling faster and more scalable learning. Our contribution provides a theoretical foundation for why stability is essential in designing efficiently parallelizable nonlinear SSMs.
@@
While this work focuses on establishing the fundamental concepts and theoretical foundations, several practical considerations arise when scaling to large systems. Notably, DEER incurs a significant memory footprint. While this issue can be alleviated through quasi-Newton methods (Gonzalez et al., 2024; Zoltowski et al., 2025), these approaches require more optimization steps to converge. Studying quasi-Newton methods in light of our theory could provide new insight into the efficacy of these methods.
Another important consideration is the choice of merit function, which is not unique. For example, one may employ a weighted norm in place of the standard Euclidean (two-)norm. By carefully designing this norm, one can potentially precondition the optimization problem, mitigating the poor conditioning often associated with Euclidean loss functions.
@@ Acknowledgments
We thank members of the Linderman Lab for helpful feedback, particularly Noah Cowan, Henry Smith, and Skyler Wu.
@@
We thank Stanford University and the Stanford Research Computing Center for providing computational resources and support. Additional computations were performed on Marlowe (Kapfer et al., 2025), Stanford University’s GPU-based Computational Instrument, supported by Stanford Data Science and Stanford Research Computing.
@@ Appendix A Brief Overview of DEER/DeepPCR
This section provides background on DEER/DeepPCR needed to support section 4 of the main text. Other options for further background on DEER are sections 2-4 of Gonzalez et al. [2024] and the corresponding blog post [Gonzalez, 2024].
We begin with a brief review of DEER/DeepPCR [Danieli et al., 2023, Lim et al., 2024, Gonzalez et al., 2024]. As mentioned in the introduction, the choice of optimizer is crucial for this procedure to outperform sequential evaluation in terms of wall clock time. Indeed, for this reason DEER uses the Gauss-Newton method (GN) to minimize the residual loss, since GN exhibits quadratic convergence rates near the optimum [Nocedal and Wright, 2006]. Recall from eq.˜3 that the $i$-th step of the DEER algorithm is,
@@
This step requires inverting the $TD\times TD$ matrix, $\mathbf{J}(\mathbf{s}^{(i)})$. Rather than explicitly inverting it, which is generally infeasible, DEER solves for the updates by running a linear time-varying recursion [Gonzalez et al., 2024]:
@@
Unlike the standard sequential rollout, this recursion can be parallelized and computed in $O(\log T)$ time using a parallel associative scan [Blelloch, 1990]. When the number of optimization steps needed for DEER to converge to the true trajectory is relatively small, DEER can yield faster overall evaluation than the sequential approach. Since Gauss–Newton converges quadratically when the initial guess is sufficiently close to the true optimum [Lim et al., 2024, Nocedal and Wright, 2006], DEER potentially only requires a tiny number of iterations to converge. Our first key result is to prove that DEER always converges globally with linear rate, and will thus always reach this basin of quadratic convergence after sufficient time.
@@
[…]
  • merit function $\mathbbR\mathcal{L}(\mathbf{s})\in\mathbb{R}$
[…]
@@ Appendix B Merit Function is PL
This section provides a proof of main text Proposition˜1. We first note that Proposition˜1 applies to optimizing any nonlinear sum of squares problem where $\mathcal{L}(\mathbf{s})=\frac{1}{2}\|\mathbf{r}(\mathbf{s})\|_{2}^{2}$, not just the $\mathbf{r}$ we consider in this paper (defined in eq.˜2).
Proposition (Proposition˜1).
@@
Substituting these expressions into the PL inequality in eq.˜8 and dropping the explicit dependence on $\mathbf{s}$ for simplicity, we obtain,
$\mathbf{r}^{\top}\mathbf{J}\mathbf{J}^{\top}\mathbf{r}\ \geq\ \mu\,\mathbf{r}^{\top}\mathbf{r}.$
@@
[…]
@@
The PL inequality or gradient dominance condition is stated differently in different texts [Nesterov and Polyak, 2006, Fazel et al., 2018, Chewi and Stromme, 2025]. We follow the presentation of Karimi et al. [2016]. Karimi et al. [2016] emphasizes that PL is often weaker than many other conditions that had been assumed in the literature to prove linear convergence rates.
@@ Appendix C Merit Function PL Constant is Controlled by Largest Lyapunov Exponent of Model
[…]
@@
[…]
@@
i.e. $\mathbf{N}(\mathbf{s})$ collects the Jacobians of the dynamics function along the first lower diagonal. Each matrix power $\mathbf{N}^{k}$ therefore collects length $k$ products along the $k$th lower diagonal. Thus, multiplication by $\mathbf{J}(\mathbf{s})^{-1}=\sum_{k=0}^{T-1}\,\mathbf{N}(\mathbf{s})^{k}$ recovers running forward a linearized form of the dynamics, which is one of the core insights of DeepPCR and DEER [Danieli et al., 2023, Lim et al., 2024].
@@
[…]
For $\rho<1$, this assumption corresponds to global contraction of the dynamics [Lohmiller and Slotine, 1998].
@@
[…]
@@
[…]
@@
We observe that as $\mathbf{s}^{(i)}$ approaches the true solution $\mathbf{s}^{*}$, the regularity conditions in eq.˜10 become increasingly reasonable. Since any successful optimization trajectory must eventually enter a neighborhood of $\mathbf{s}^{*}$, it is natural to expect these conditions to hold there. In fact, rather than requiring the regularity conditions over all of state space or along the entire optimization trajectory, one could alternatively assume that they hold within a neighborhood of $\mathbf{s}^{*}$, and prove a corresponding version of Theorem˜2.
@@
A consequence of Theorem˜6 is that if the system is unpredictable, then there exists a finite ball around $\mathbf{s}^{*}$ where the conditioning of the merit function landscape is provably bad.
@@
In our proof of Theorem˜2, we proved upper and lower bounds for $\sigma_{\min}(\mathbf{J}(\mathbf{s}))$ that depended on the sequence length $T$. We can also prove upper and lower bounds for $\sigma_{\max}(\mathbf{J}(\mathbf{s}))$, but these do not depend on the sequence length.
@@
[…]

### v2

@@ 6 Conclusion
The DeepPCR algorithm was introduced by Danieli et al. [2], who investigated its convergence rates empirically but not theoretically. Concurrently, Lim et al. [1] proposed an essentially identical method under the name DEER, proving local quadratic convergence for Gauss–Newton under standard assumptions but leaving the question of global convergence unresolved. Gonzalez et al. [3] proved the global convergence of DEER and other variants, though only with worst-case bounds of $T$ optimization steps. None of these prior works addressed the relationship between system dynamics and conditioning, or established global linear convergence rates.
@@
Global convergence rates for Gauss-Newton are rare, despite the breadth of optimization literature [19, 28, 33]. Theorem 4 establishes global convergence with linear rate for Gauss-Newton by leveraging our specific problem structure.
@@
Fifty years ago, Hyafil and Kung [34] and Kung [35] showed that linear recursions enjoy speedups from parallel processors while nonlinear recursions of rational functions with degree larger than one cannot. These prescient works set the stage for our more general findings, which explicitly link the dynamical properties of the recursion to its parallelizability. Parallel-in-time methods for continuous systems also have a long history [36, 37], with Chartier and Philippe [38] showing that dissipative systems can be parallelized using multiple shooting. Furthermore, Danieli and MacLachlan [39] and De Sterck et al. [40] study the CFL number for determining the usefulness of multigrid systems. Connecting this work with our paper is an interesting direction for future research.
@@
More recently, several works have parallelized diffusion models via fixed-point iteration [2, 11, 41, 42], again with $T$-step worst-case guarantees. Anari et al. [43] proved $\log(T)$ rates for a particular dynamical system (Langevin dynamics) and a particular fixed-point iteration (Picard iteration). Crucially, prior work has not focused on the merit function, which we can define for any discrete-time dynamical system and optimizer.
@@
To our knowledge, no prior work connects the LLE of a dynamical system to the conditioning of the corresponding optimization landscape, as established in ˜2. In particular, we showed that systems with high unpredictability yield poorly conditioned (i.e., flat) merit functions, linking dynamical instability to optimization difficulty in a geometrically appealing way.
@@
The centrality of parallel sequence modeling architectures like transformers [44], deep SSMs [45, 21, 46], and linear RNNs [47] in modern machine learning underscores the need for our theoretical work. Merrill et al. [48] explored the question of parallelizability through the lens of circuit complexity, analyzing when deep-learning models can solve structured tasks in constant depth. Their focus complements ours, and suggests an opportunity for synthesis in future work.
@@
Nonlinear state space models are ubiquitous. Our results offer two key contributions.
@@
First, they provide a principled way to determine, a priori, whether optimization-based parallelization of a given model is practical. In many robotic or control systems, particularly ones that are strongly dissipative, this insight can enable orders-of-magnitude speed-ups on GPUs [49, 50, 51, 52, 53, 54, 55, 56].
@@
In concurrent work, Zoltowski et al. [18] developed and leveraged quasi-Newton methods to parallelize Markov Chain Monte Carlo over the sequence length, attaining order-of-magnitude speed-ups. These speed-ups were made possible by the fast convergence of the quasi-Newton methods in the settings considered. Suggestively, there is an abundance of research studying the contractivity of MCMC in different settings [57, 58, 59]. In fact, the empirical results in this paper showing that Langevin dynamics can have negative LLE (cf. Figure˜3) are suggestive that the Metropolis-adjusted Langevin algorithm (MALA), a workhorse of MCMC, may also be predictable in settings of interest. A precise characterization of what makes an MCMC algorithm and target distribution contractive (i.e. predictable, in the language of our paper) would provide useful guidance for when one should aim to parallelize MCMC over the sequence length. Thus, providing precise theoretical justification for parallelizing MCMC over the sequence length is an exciting avenue for future work.
@@
Second, our results have direct implications for system design. When constructing nonlinear dynamical systems in machine learning—such as novel RNNs—parallelization benefits are maximized when the system is made predictable. Given the large body of work on training stable RNNs [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], many effective techniques already exist for enforcing stability or predictability during training. A common approach is to parameterize the model’s weights so that the model is always stable. For example, we provide a simple parameterization of a contractive SSM in Appendix˜I.
@@
Notably, the concurrent work of Farsang et al. [71] develops a nonlinear SSM and trains it in parallel using quasi-DEER [3]. Farsang et al. [71] explicitly parameterizes their LrcSSM to be contractive (see their Appendix A.1). While they do not identify it as such, this contractivity is precisely what enables DEER to converge quickly throughout training. Ensuring a negative largest Lyapunov exponent through parameterization guarantees parallelizability for the entire training process, enabling faster and more scalable learning. Our contribution provides a theoretical foundation for why stability is essential in designing efficiently parallelizable nonlinear SSMs.
@@
While this work focuses on establishing the fundamental concepts and theoretical foundations, several practical considerations arise for scaling to large systems. Notably, DEER incurs a significant memory footprint. While this issue can be alleviated through quasi-Newton methods [3, 18], these approaches require more optimization steps to converge. Studying quasi-Newton methods with our theory could provide new insight into the efficacy of these methods.
@@ Acknowledgments
We thank members of the Linderman Lab for helpful feedback, particularly Noah Cowan, Henry Smith, Skyler Wu, and Etaash Katiyar. We thank Federico Danieli for helpful discussions.
@@
We thank Stanford University and the Stanford Research Computing Center for providing computational resources and support. Additional computations were performed on Marlowe [72], Stanford University’s GPU-based Computational Instrument, supported by Stanford Data Science and Stanford Research Computing.
@@ Appendix A Brief Overview of DEER/DeepPCR
This section provides background on DEER/DeepPCR needed to support section 4 of the main text. Other options for further background on DEER are sections 2-4 of Gonzalez et al. [3] and the corresponding blog post [73].
We begin with a brief review of DEER/DeepPCR [2, 1, 3]. As mentioned in the introduction, the choice of optimizer is crucial for this procedure to outperform sequential evaluation in terms of wall clock time. Indeed, for this reason DEER uses the Gauss-Newton method (GN) to minimize the residual loss, since GN exhibits quadratic convergence rates near the optimum [19]. Recall from eq.˜3 that the $i$-th step of the DEER algorithm is,
@@
This step requires inverting the $TD\times TD$ matrix, $\mathbf{J}(\mathbf{s}^{(i)})$. Rather than explicitly inverting it, which is generally infeasible, DEER solves for the updates by running a linear time-varying recursion [3]:
@@
Unlike the standard sequential rollout, this recursion can be parallelized and computed in $O(\log T)$ time using a parallel associative scan [5]. When the number of optimization steps needed for DEER to converge to the true trajectory is relatively small, DEER can yield faster overall evaluation than the sequential approach. Since Gauss–Newton converges quadratically when the initial guess is sufficiently close to the true optimum [1, 19], DEER potentially only requires a tiny number of iterations to converge. Our first key result is to prove that DEER always converges globally with linear rate, and will thus always reach this basin of quadratic convergence after sufficient time.
@@
  • residual $\mathbf{r}(\mathbf{s})\in\mathbb{R}^{TD}$
  • Jacobian $\mathbf{J}(\mathbf{s})\in\mathbb{R}^{TD\times TD}$
  • merit function $\mathcal{L}(\mathbf{s})\in\mathbb{R}$
are functions of the current guess for the trajectory $\mathbf{s}=\mathrm{vec}(s_{1},\ldots,s_{T})\in\mathbb{R}^{TD}$. As much as possible, we try to emphasize the dependence on the current guess for the trajectory, but sometimes we will drop the dependence for notational compactness.
@@ Appendix B Merit Function is PL
This section provides a proof of main text ˜1. We first note that ˜1 applies to optimizing any nonlinear sum of squares problem where $\mathcal{L}(\mathbf{s})=\frac{1}{2}\|\mathbf{r}(\mathbf{s})\|_{2}^{2}$, not just the $\mathbf{r}$ we consider in this paper (defined in eq.˜2).
Proposition (˜1).
@@
Substituting these expressions into the PL inequality in eq.˜8 we obtain
$\mathbf{r}^{\top}\,\mathbf{J}(\mathbf{s})\,\mathbf{J}(\mathbf{s})^{\top}\,\mathbf{r}\ \geq\ \mu\,\mathbf{r}^{\top}\mathbf{r}.$
@@
To be precise, we must have $\mu>0$ for $\mathcal{L}$ to satisfy the definition of PL. Therefore, a condition that must apply for $\mathcal{L}$ to be PL is that we must have $\inf_{\mathbf{s}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$. We note that the proof strategy of ˜2 ensures that $\inf_{\mathbf{s}\in\mathbb{R}^{TD}}\sigma_{\min}\left(\mathbf{J}(\mathbf{s})\right)>0$ if we assume eq.˜21, which holds for dynamical systems that are globally contracting.
By the chain rule, eq.˜21 also holds for functions of the form $f(s)=\phi(Ws)$, where $W\in\mathbb{R}^{D\times D}$ and $\phi$ is a scalar function with bounded derivative that is applied elementwise. In particular, such a function $\phi(Ws)$ satisfies eq.˜21 whether or not it is globally contracting. This function class is extremely common in deep learning (nonlinearities with bounded derivatives include $\tanh$, the logistic function and $\mathrm{ReLU}$).
In our statement and proof of ˜1, we deliberately do not specify the set over which we take the infimum. The result is true regardless of what this set is taken to be. The largest such set would be $\mathbb{R}^{TD}$, but other sets that could be of interest are the optimization trajectory $\{\mathbf{s}^{(i)},i\in\mathbb{N}\}$, or alternatively a neighborhood of the solution $\mathbf{s}^{*}$. We discuss further in Appendix˜C.
@@
The PL inequality or gradient dominance condition is stated differently in different texts [27, 26, 74]. We follow the presentation of Karimi et al. [25]. Karimi et al. [25] emphasizes that PL is often weaker than many other conditions that had been assumed in the literature to prove linear convergence rates.
@@ Appendix C Merit Function PL Constant is Controlled by Largest Lyapunov Exponent of Model
We pick two unit vectors $u$ and $v$, both in $\mathbb{R}^{TD}$, that are zero everywhere other than where they need to be to pull out the bottom-left block of $\mathbf{J}^{-1}$ (i.e., the only non-zero block in $\mathbf{N}^{T-1}$, which is equal to $J_{T}J_{T-1}\ldots J_{2}$). Doing so, we get
@@
where $\tilde{u}$ and $\tilde{v}$ are unit vectors in $\mathbb{R}^{D}$, and are equal to the nonzero entries of $u$ and $v$.
@@
i.e. $\mathbf{N}(\mathbf{s})$ collects the Jacobians of the dynamics function along the first lower diagonal. Each matrix power $\mathbf{N}^{k}$ therefore collects length $k$ products along the $k$th lower diagonal. Thus, multiplication by $\mathbf{J}(\mathbf{s})^{-1}=\sum_{k=0}^{T-1}\,\mathbf{N}(\mathbf{s})^{k}$ recovers running forward a linearized form of the dynamics, which is one of the core insights of DeepPCR and DEER [2, 1].
@@
$\sup_{s\in\mathbb{R}^{D}}\|J(s)\|_{2}\leq\rho.$ (21)
For $\rho<1$, this assumption corresponds to global contraction of the dynamics [7].
@@
$\frac{1}{a}\cdot\frac{\rho-1}{\rho^{T}-1}\leq\sqrt{\inf_{\mathbf{s}\in\mathbb{R}^{TD}}\sigma^{2}_{\min}(\mathbf{J}(\mathbf{s}))}.$
@@
The LLE regularity conditions in eq.˜10 highlight the more natural “average case” behavior experienced along actual trajectories $\mathbf{s}\in\mathbb{R}^{TD}$. This “average case” behavior is highlighted, for example, by our experiments with the two-well system (cf. Section 5 and Appendix J.4), where even though a global upper bound on $\|J_{t}(s_{t})\|_{2}$ over all of state space would be greater than $1$ (i.e., there are unstable regions of state space), we observe fast convergence of DEER because the system as a whole has negative LLE (its trajectories are stable on average).
@@
We observe that as $\mathbf{s}^{(i)}$ approaches the true solution $\mathbf{s}^{*}$, the regularity conditions in eq.˜10 become increasingly reasonable. Since any successful optimization trajectory must eventually enter a neighborhood of $\mathbf{s}^{*}$, it is natural to expect these conditions to hold there. In fact, rather than requiring the regularity conditions over all of state space or along the entire optimization trajectory, one could alternatively assume that they hold within a neighborhood of $\mathbf{s}^{*}$, and prove a corresponding version of ˜2.
@@
A consequence of ˜6 is that if the system is unpredictable, then there exists a finite ball around $\mathbf{s}^{*}$ where the conditioning of the merit function landscape is provably bad.
@@
In our proof of ˜2, we proved upper and lower bounds for $\sigma_{\min}(\mathbf{J}(\mathbf{s}))$ that depended on the sequence length $T$. We can also prove upper and lower bounds for $\sigma_{\max}(\mathbf{J}(\mathbf{s}))$, but these do not depend on the sequence length.
@@
$\sup_{\mathbf{s}\in\mathbb{R}^{TD}}\|\mathbf{N}(\mathbf{s})\|_{2}$ $=\sup_{s\in\mathbb{R}^{D}}\|J(s)\|$
Elaborating, for a particular choice of trajectory $\mathbf{s}\in\mathbb{R}^{TD}$, $\|\mathbf{N}(\mathbf{s})\|_{2}$ is controlled by the maximum spectral norm of the Jacobians $J_{t}(s_{t})$ along this trajectory. Analogously, $\sup_{\mathbf{s}\in\mathbb{R}^{TD}}\|\mathbf{N}(\mathbf{s})\|_{2}$—i.e., the supremum of the spectral norm of $\mathbf{N}(\mathbf{s})$ over all possible trajectories $\mathbf{s}\in\mathbb{R}^{TD}$, i.e. the optimization space—is upper bounded by $\sup_{s\in\mathbb{R}^{D}}\|J(s)\|_{2}$, i.e. the supremum of the spectral norm of the system Jacobians over the state space $\mathbb{R}^{D}$.